SUBMILLIMETER-WAVE ROTATIONAL SPECTROSCOPY OF H_{2}F^{+}

\underline{R. FUJIMORI}, K. KAWAGUCHI, Department of Chemistry, Faculty of Science, Okayama University, 3-1-1, Tsushima-Naka, Okayama 700-8530, Japan; T. AMANO, Department of Chemistry and Department of Physics and Astronomy, University of Waterloo, 200 University Avenue West, Waterloo, ON N2L 3G1, Canada.

Five pure rotational transitions of H_{2}F^{+} were observed in the 473-774 GHz range with a backward-wave oscillator based submillimeter-wave spectrometer^{a}. The H_{2}F^{+} ion was generated in an extended negative glow discharge in a gas mixture of hydrogen fluoride generated by heating potassium hydrogen fluoride (HF_{2}K) granular powder at 150^{∘}C-160^{∘}C and hydrogen in an argon buffer. A simultaneous analysis of the rotational lines with 120 combination differences for the ground state derived from the infrared spectra obtained by Schäfer and Saykally^{b} and Fujimori et al.^{c} was carried out to determine the precise molecular constants for the ground state. The rotational transition frequencies that lie below 2 THz were calculated, together with their estimated uncertainties, to facilitate future astronomical identifications. Recently H_{2}Cl^{+} was detected in NGC 6334I and Sgr B2 with the Heterodyne Instrument for Far-Infrared (HIFI) on board the Herschel Space Observatory^{d}, and HF was also detected in a wide variety of interstellar clouds with the same facility^{e}. The proton affinity of HF is smaller than those of N_{2} and CO, so the abundance of H_{2}F^{+} is likely to be low in dense molecular clouds. We will discuss abundances of H_{2}F^{+} in diffuse molecular clouds, considering various chemical reaction rates.

aT. Amano and T. Hirao, J. Mol. Spectrosc., 233, 7 (2005)

bE. Schäfer and R. J. Saykally, J. Chem. Phys., 81, 4189 (1984)

cR. Fujimori, Y. Hirata, K. Kawaguchi, and I. Morino, RE02, The 65^{th} OSU International Symposium on Molecular Spectroscopy (2010)(Columbus, OH)

dD. C. Lis et al., Astron. Astrophys., 521, L9 (2010)

eD. A. Neufeld et al., Astron. Astrophys., 518, L108 (2010); T. G. Phillips et al., Astron. Astrophys., 518, L109 (2010); P.

Sonnentrucker et al., Astronon. Astrophys., 521, L12 (2010)